Find the Laplace Transform of the following functions:

i)ft=e-tsinh 3t

                                                       ii)ft=t9e5t


1
Expert's answer
2021-07-30T04:18:01-0400

i

"\mathrm{Apply\:transform\:rule:\quad if\:}L\left\{f\left(t\right)\right\}=F\left(s\right)\mathrm{\:then\:}\:L\left\{e^{at}f\left(t\right)\right\}=F\left(s-a\right)\\
\mathrm{For\:}e^{-t}\sinh \left(3t\right):\quad f\left(t\right)=\sinh \left(3t\right),\:\quad a=-1\\
=L\left\{\sinh \left(3t\right)\right\}\left(s+1\right)\\
=\frac{3}{s^2-9}\\
=\frac{3}{\left(s+1\right)^2-9}"

ii

"\mathrm{Use\:Laplace\:Transform\:table}:\quad \:L\left\{t^kf\left(t\right)\right\}=\left(-1\right)^k\frac{d^k}{ds^k}\left(L\left\{f\left(t\right)\right\}\right)\\
\mathrm{For\:}t^9e^{5t}:\quad f\left(t\right)=e^{5t},\:\quad \:k=9\\
=\left(-1\right)^9\frac{d^9}{ds^9}\left(L\left\{e^{5t}\right\}\right)\\
L({e^{5t}})=\frac{1}{s-5}\\
\frac{d^9}{ds^9}(\frac{1}{s-5})=-\frac{362880}{\left(s-5\right)^{10}}\\
=\left(-1\right)^9\left(-\frac{362880}{\left(s-5\right)^{10}}\right)\\
=\frac{362880}{\left(s-5\right)^{10}}"